OpenStudy (swaugh):

MEDAL + FAN AWAITINGGGGG What is the slope of a line that passes through the coordinates two comma twelve and the coordinates negative one comma three? Hint: Think rise over run. -3 - 1/3 1/3 3

OpenStudy (abmon98):

m(slope)=y2-y1/x2-x1= plug in your givens and your done.

OpenStudy (anonymous):

use the slope formula\[\huge{m=\frac{ y_2-y_1 }{ x_2-x_1 }}\]\[\huge{Let~say(2,12)=(x_1,y_2)}\]\[\huge{Let~say(-1,3)=(x_2,y_2)}\]Then,plug in the value into the slope formula\[\huge{m=\frac{ y_2-y_1 }{ x_2-x_1 }}\]\[\huge{m=\frac{ 3-12 }{ -1-2 }}\]\[\huge{m=\frac{ -9 }{ -3 }}\]Then,try to reduce the result. We notice that -9 and-3 can be divide by -3.\[\huge{m=\frac{ 3\times-3 }{ 1\times-3 }}\]\[\huge{m=\frac{ 3\times\cancel{-3} }{ 1\times\cancel{-3} }}\]\[\huge{m=?}\]solve it @swaugh

OpenStudy (swaugh):

is it 3

OpenStudy (anonymous):

correct!